OpenStudy (anonymous):

Does anyone know an algebraic expression that gives the area of a circle with a circumference ?

OpenStudy (anonymous):

circumfrance = pi * d

OpenStudy (anonymous):

area = p * r^2

OpenStudy (anonymous):

ok so A= pi*r^2*d ?

OpenStudy (anonymous):

no i dont' think that there is one for both. you have to use diff. values for each one

OpenStudy (anonymous):

ok thanks!

OpenStudy (amistre64):

Area of a circle = Circumference times r /2 C = 2pi(r) A = (2pi(r))r/2 A= pi(r^2)

OpenStudy (amistre64):

Actually the Area of a sector is equal to its degree in radians times the square of the radius divided by 2. The radian measure of a circle is (2pi). 2pi(r^2)/2 = pi(r^2)

OpenStudy (anonymous):

The question for this one says, Enter here an algebraic expression that gives the area of a circle with a circumference . Hint: The circumference of a circle with radius is , and its area is

OpenStudy (amistre64):

Algebraic expressions: Circumference = 2pi(r) Area = pi(r^2) When you know the radius, you can solve :)

OpenStudy (anonymous):

isnt the radius 1/2

OpenStudy (anonymous):

radius = d/2

OpenStudy (amistre64):

the radius depends on the size of the circle. And it is half of the width.

OpenStudy (anonymous):

yeah so diameter/ 2